# Neutron and $\bm{\gamma}$-ray Discrimination by a Pressurized Helium-4 Based Scintillation Detector

### v2

@@ Abstract
Pressurized helium-4 based fast neutron scintillation detector offers an useful alternative to organic liquid-based scintillator due to its relatively low response to the $\gamma$-rays compared to the latter type of scintillator. In the present work, we have investigated the capabilities of a pressurized 4He (PHe) detector for the detection of fast neutrons in a mixed radiation field where both the neutrons and the $\gamma$-rays are present. Discrimination between neutrons and $\gamma$-rays is achieved by using fast-slow charge integration method. We have also conducted systematic studies of the attenuation of fast neutrons and $\gamma$-rays by high-density polyethylene (HDPE). These studies are further corroborated by simulation analyses conducted using GEANT4, which show qualitative agreement with the experimental results. Additionally, the simulation provides detailed insights into the interactions of the radiation quanta with the PHe detector. Estimates of the scintillation signal yield are made based on our GEANT4 simulation results by considering the scintillation mechanism in the PHe gas.
keywords:
Pressurized helium-4 detector , neutron - gamma discrimination , neutron spectroscopy , neutron attenuation , GEANT4 simulation
[…]
[affil1] organization=High Energy Nuclear and Particle Physics Division, Saha Institute of Nuclear Physics - a CI of Homi Bhabha National Institute, addressline=Block AF, Sector I, city=Bidhannagar, Kolkata, postcode=700064, state=, country=India
[…]
[affil3] organization=Applied Nuclear Physics Division, Saha Institute of Nuclear Physics - a CI of Homi Bhabha National Institute, addressline=Block AF, Sector I, city=Bidhannagar, Kolkata, postcode=700064, state=, country=India
[…]
[affil2] organization=Department of Physics and Astronomy, Purdue University, city=West Lafayette, postcode=47907, state=IN, country=USA
@@ 1 Introduction
Neutrons as radiation quanta, are found in nature because of their release in various forms of nuclear reactions, most common being the nuclear fission and the $(\alpha,n)$ reactions caused by the natural radioactivity of the remnants of Uranium-Thorium (U-Th) decay chain. Free neutrons are, however, unstable against $\beta$-decay with about 15 minutes of half life. In spite of this fact, free neutrons pass and penetrate through matter since they are charge neutral and deposits partial or full energy through hadronic interactions. These interactions include capture into nuclei and elastic scattering, resulting in nuclear recoil of the stopping media. Such interactions take place over a very short time scale compared to the half-life of free neutrons, which make it possible to detect them in real life.
@@
Liquid helium[3] has been investigated as scintillator for neutron detection more than 60 years back. However, the scintillation properties of PHe gas have been examined much later for successful implementation as fast neutron detector[4]. The major advantage of PHe gas as scintillator is its relatively weak response to $\beta$-particles and $\gamma$-rays, because the density of available electrons in helium is much less than in standard scintillators, organic and inorganic. On the other hand, neutrons cause nuclear recoil of the helium-4 nuclei inside the pressurized gas, resulting in multiple processes of ionization and other interactions, leading to scintillation through transitions from the singlet excimer states or the triplet excimer states[5]. Both the transitions lead to emission in the extreme ultraviolet (EUV) region of the electromagnetic spectrum, with wavelength of maximum emission around 80 nm. A wavelength shifting (WLS) compound is used to make the EUV scintillation light output readable by photomultipliers (PMTs) or silicon photomultipliers (SiPMs)[6, 7].
@@
In recent times, PHe detectors are made commercially available by Arktis Radiation Detectors Ltd, Switzerland[8]. Front-end electronics and data acquisition system is provided as a package for measuring the fast and the thermal neutron fluxes after pulse shape discrimination using Time-over-Threshold (ToT) technique to discriminate between $\gamma$-rays, fast neutrons and the thermal neutrons. The inner wall of the detector is coated with a Lithium-based compound to make the detector sensitive to the thermal neutrons. These neutrons undergo capture by Lithium to produce energetic $\alpha$-particles which result in scintillation inside the detector volume. Corresponding ToT signals are found to be larger than those produced by the fast neutrons. Some technical details about the variant detector model S670, which is not sensitive to the thermal neutrons, can be found elsewhere[9].
Aim of this work is the following: 1) discrimination of fast neutrons from the $\gamma$-rays and electrons using fast-slow charge integration method; 2) qualitative study of attenuation of the fast neutrons and $\gamma$-rays by high density polyethylene (HDPE); 3) estimation of threshold for neutron and $\gamma$-ray discrimination; 4) carry out a detailed simulation to understand the interaction of the neutrons and $\gamma$-rays with the PHe media; 5) follow the systematics of energy transfer from the neutrons and the $\gamma$-rays to 4He (nuclear recoil) and electrons (electron recoil) respectively; and 6) to estimate the resulting scintillation signals following excitation and charge transfer collision processes in the PHe.
@@ 2.1 Instrumentation
A photograph of the experimental set up is shown in the Fig. 1(a). Arktis-made PHe detector and a radioactive source (252Cf source inside a PTFE capsule), placed in front of the detector, are displayed in the photograph. The active medium of the detector is 4He gas at a pressure of 150 - 180 bar, enclosed inside a stainless steel tube of 60 mm inner diameter and 600 mm active length. Details about the detector, related electronics and their working principles are given in many references available as published articles[6, 10]. The detector is packaged with SiPM arrays as photon readout located inside the pressurized detector volume. The detector is segmented into three parts along the length of the cylindrical tube. Photons from each segment are read out by an array of 8 SiPMs. Signals from two SiPMs are summed and fed into each output, so that 4 output signal pulses are generated from each segment.
@@
For our work, the digital electronics readout system was replaced by a multichannel analog circuit (provided by the manufacturer) for analog signal processing using conventional electronics and a 12-bit current integrating analog to digital converter or QDC (16 channel Phillips Scientific charge ADC, Model 7187). The pulses from the SiPMs are fed to the analog circuit, which consists of a preamplifier and a shaping amplifier with baseline restorer for each SiPM channel. Traces of the output pulses from the amplifier are shown in the Fig. 1(b). The output signals were inverted using high speed pulse transformer (ALT 4532M) to match with the polarity and pulse timing requirements. Out of the four output signals from a segment, one signal was fed to a low threshold discriminator to generate the logic gates for the QDCs required for pulse shape discrimination (PSD) between neutrons and $\gamma$-rays or electrons. The PSD parameter ($P$) is defined as: $P=NQ_{S}/{(Q_{S}+Q_{L})}$ where, $Q_{S}$ and $Q_{L}$ are the charge contents of the pulse within the duration of the short gate and the long gate respectively and $N$ is a scaling factor to convert the ratios to suitable integer values for the plots.
@@ 2.2 Measurements with Radioactive sources
The systematic studies were done using different neutron and $\gamma$-ray sources placed at a certain distance from the detector as shown in the photograph (see Fig. 1(a)). Provision for placement of different absorber materials in between for systematic studies was also made. An unmoderated 252Cf spontaneous fission source (half-life = 2.645 years, spontaneous fission branching ratio = 3.09%[11]) was used for simultaneous detection of neutrons and $\gamma$-rays. The source emits fast neutrons with average energy of 2.12 MeV. The source, with a strength of $\sim 1-5\,{\rm kBq}$, was sealed inside a cylindrical PTFE capsule having an opening on one of the flat faces. The aperture diameter was 2 mm, which was sealed with a 100 $\mu$m thick polyethylene terepthalate (PET) window.
A typical scatter plot of the PSD parameter $P$ vs. $Q_{L}$ is shown in Fig. 2 for the 252Cf source placed near the detector. Two distinct bands can be seen in the scatter plot
Figure 2: Scatter plot of the PSD parameter $P$ as function of $Q_{L}$ obtained by exposure of 252Cf source. Inset shows the gate configurations for the short and the long gates.
obtained after 5 hours of exposure to the 252Cf source. The relative configuration of the gates, optimized by adjusting gate widths and delays by looking at $P$, is also shown in the Fig. 2. Optimum delay between the threshold for the up-swing of the pulse and the trigger point of the short gate was found to be around 300 nanosecond for achieving good discrimination.
A 137Cs monoenergetic (662 keV) $\gamma$-ray source, with strength $\sim 30\,{\rm kBq}$, was used for the experiment to identify the $\gamma$-band as distinct from the neutron band. The scatter plot, obtained after exposing the detector with the $\gamma$-ray source, is displayed in the Fig. 3, which shows a single band ($\gamma$-band) as expected.
@@
From the scatter plot of Fig. 2, the neutron and the $\gamma$-bands are found to merge at low $Q_{L}$ values, which qualitatively indicates the low energy threshold for discrimination. A comparison of projection of the neutron and the $\gamma$-bands on the $Q_{L}$ axis is shown in the Fig. 4. It reveals that the detector response to the $\gamma$-rays, mediated mostly through electron recoil, is quite low compared to that for the neutrons having energies of the same order. It can be seen from the plots that the nuclear recoil spectrum due to the neutrons terminate abruptly at higher channels ($\gtrsim 250$), which is due to the saturation of the pulses of the SiPM signal amplifiers provided by the manufacturer.
@@
2.3 Measurements with 241Am X-ray source
A 241Am X-ray source, embedded in a glass matrix and sealed inside a cylindrical stainless steel capsule, was also studied with the detector. The source, having a strength of $\sim 100\,{\rm MBq}$, is originally intended for energy calibration of X-ray detectors. This source was used in the experiment to determine the response of the detector to low energy photons so that a low energy cut-off for electron recoil could be established. The scatter plot of $P$ vs. $Q_{L}$, obtained after 5 hours of exposure, is shown in the Fig. 5. A clear neutron band, besides the $\gamma$-band, was observed to appear. This was a bit surprising due to the reason that 1) the source was used so far for energy and relative efficiency calibration of the X-ray detectors, and 2) it was never monitored with neutron detectors or dosimeters. Surprisingly, the same exposure time (5 hours) for the 241Am and the 252Cf sources to the PHe detector was used.
In order to confirm that the band is due to the neutrons being detected, attenuation of the band population by high density polyethylene (HDPE) was studied. For this purpose, 4 layers of HDPE, each being 25 mm thick with a total thickness of 100 mm, were placed between the 241Am source and the PHe detector to attenuate the neutrons. The scatter plot ($P$ vs. $Q_{L}$), obtained over the same exposure time, is shown in the Fig. 5.
@@ 2.3 Measurements with 241Am X-ray source
It clearly reveals attenuation of the neutron band. The plots, obtained by gating on the neutron band and projection taken along the $Q_{L}$ axis (see Fig. 6), provide a comparison of the neutron attenuation by the HDPE layers. Low energy neutrons are expected to be attenuated more than those of higher energies, however from the plot, the attenuation factor appears to be fairly independent of energy. This is due to the moderation of the energetic neutrons that enhances counts at low energy and suppresses at higher energy. Similar projection spectrum for the $\gamma$-band is shown in the Fig. 7, where attenuation of $\gamma$-rays is observed. We also have observed build-up of counts in the high energy tail (larger $Q_{L}$). This was investigated in detail through simulation in Sec. 3.4.
@@
Figure 7: Same as Fig. 6 except that the banana gate is set on the $\gamma$-band. (a) projected spectra with and without the HDPE layers between the source and the detector. See Sec. 3.4 for the calibration procedure applied in the plot.(b) Comparison of the projected $\gamma$-spectra recorded after the neutrons travel through different numbers of the HDPE layers. The $x$-axis scale on the bottom displays the electron equivalent energy loss $E_{dep}^{ee}$ of neutron. See Sec. 3.3 for its definition.
3 Results and analysis
3.1 Simulation
A detailed simulation of response systematic of the detector to the relevant radiation quanta (neutrons, $\gamma$-rays) were done to understand the nuclear-recoil events (NR) due to the neutrons and the electron-recoil (ER) events due to the $\gamma$-rays, and to compare them qualitatively with the results obtained from the experiment. It is evident from the previous section that the scope of simulation are 5-fold: 1) emission of relevant radiation quanta from the radioactive source used, particularly the 241Am source; 2) response of the detector to the neutrons, more specifically to the fast neutrons; 3) attenuation of the fast neutrons by suitable plastic absorbers (HDPE); 4) response of the detector to the $\gamma$-rays emitted from the radiation sources and also resulting from neutron absorption in HDPE; 5) discrimination between the neutrons and the $\gamma$-rays using signal analysis.
Simulations were carried out by the GEANT4 (G4) simulation toolkit [12], version 10.7.4. G4 provides numerous handles to tune the simulation parameters, allowing users to focus on specific aspects of interest at a time. A custom physics-list is utilized in our simulation, which is based on the one developed by Mendoza et al.[13]. We will briefly outline the G4 packages used in this physics-list to model various interactions. The G4RadioactiveDecayPhysics package is used for modeling of the radioactive decays, relevant for the 241Am source. It uses the ENSDF database [14] for the various decay parameters including the energy levels of the daughter nuclei. The production of neutrons from ($\alpha,n$) reaction is modeled using the G4ParticleHP package, which is capable of utilizing the ENDF-6 formatted data libraries[15]. The database used for this purpose is the JENDL/AN-2005 dataset [16]. The QGSP_BIC_HP model is used for the hadronic interactions and uses the G4 Neutron Data Library (G4NDL) to implement low energy neutron interactions with high-precision. The EM interactions are modeled with the G4EmStandardPhysics_option4 package.
The geometry to obtain the particle spectra consists of a cylindrical stainless steel source capsule of 6 mm diameter $\times$ 9 mm length, with the 241Am source, embedded in a pyrex glass matrix of 3 mm diameter $\times$ 6 mm thickness, placed inside and sealed with a stainless steel cover. The source capsule is surrounded by a spherical dummy detector, which is made of air. This is placed to track the particles that are being emitted from the source.
241Am (432 years half-life) is used as the primary $\alpha$-emitter nucleus embedded within the glass matrix. It decays to 237Np by emitting $\alpha$-particles at 5.486 MeV ($\sim 85\%$ decay branch), 5.443 MeV ($\sim 13\%$ decay branch), and the rest at other energies. Np237 daughter nucleus has a half life of approximately 4 orders of magnitude more than that of 241Am. Therefore, the source is approximated as an alpha emitter whose energy is sampled from the spectrum as obtained from G4. These $\alpha$-particles are allowed to penetrate isotropically through the base material of the source capsule, where they undergo ($\alpha,n$) reaction with the constituent nuclei. Based on the cross-section data available from the database, respective elemental compositions of the materials and the relevant energies of the $\alpha$-particles, we find that the dominant reaction producing neutrons are: 11B($\alpha,n$)14N ($\sim 69\%$) and 23Na($\alpha$, n)26Al($\sim 11\%$), rest are being produced by various other reactions.
Figure 8: Simulated energy spectrum of the neutrons arising from the 241Am source capsule.
3.2 Neutron Response Systematics by G4
The geometry of the PHe detector is mentioned in Sec. 2.1. For this study, the source is placed at a fixed distance from the PHe detector to accommodate the neutron absorbers (HDPE) in between. The HDPE absorber slabs, each 25 mm thick, are cumulatively placed to record the resulting neutron energy spectral distributions as seen by the PHe detector.

### v3

# Neutron and $\boldsymbol{\gamma}$-ray Discrimination by a Pressurized Helium-4 Based Scintillation Detector

@@ Abstract
Pressurized helium-4 (PHe) based fast neutron scintillation detector offers an useful alternative to organic liquid-based scintillator due to its relatively low response to the $\gamma$-rays compared to the latter type of scintillator. In the present work, we have investigated the capabilities of a PHe detector for the detection of fast neutrons in a mixed radiation field where both the neutrons and the $\gamma$-rays are present. Discrimination between neutrons and $\gamma$-rays is achieved by using fast-slow charge integration method. We have also conducted systematic studies of the attenuation of fast neutrons and $\gamma$-rays by high-density polyethylene (HDPE). Additionally, the simulation analyses, conducted using GEANT4, provide detailed insights into the interactions of the radiation quanta with the PHe detector.
@@ 1 Introduction
Neutrons as radiation quanta, are found in nature because of their release in various forms of nuclear reactions, most common being the nuclear fission and the $(\alpha,n)$ reactions caused by the natural radioactivity of the remnants of Uranium-Thorium (U-Th) decay chain. Free neutrons are, however, unstable against $\beta$-decay with about 15 minutes of half-life. In spite of this fact, free neutrons pass and penetrate through matter since they are charge neutral and deposits partial or full energy through hadronic interactions. These interactions include capture into nuclei and elastic scattering, resulting in nuclear recoil of the stopping media. Such interactions take place over a very short time scale compared to the half-life of free neutrons, which make it possible to detect them in real life.
@@
Liquid Helium[3] has been investigated as scintillator for neutron detection more than 60 years back. However, the scintillation properties of PHe gas have been examined much later for successful implementation as fast neutron detector[4]. The major advantage of liquid Helium-4 and PHe gas as scintillators is relatively weak response to $\beta$-particles and $\gamma$-rays, because the density of available electrons in Helium is much less than in standard scintillators, organic and inorganic. On the other hand, neutrons cause nuclear recoil of the Helium-4 nuclei inside the pressurized gas, resulting in multiple processes of ionization and other interactions, leading to scintillation through transitions from the singlet excimer states or the triplet excimer states[5]. Both the transitions lead to emission in the extreme ultrared (EUV) region of the electromagnetic spectrum, with wavelength of maximum emission around 80 nm. A wavelength shifting (WLS) compound is used to make the EUV scintillation light output readable by photomultipliers (PMTs) or silicon photomultipliers (SiPMs)[6, 7].
@@
In recent times, PHe detectors are made commercially available by Arktis Radiation Detectors Ltd., Switzerland[8]. Front-end electronics and data acquisition system is provided as a package for measuring the fast and the thermal neutron fluxes after pulse shape discrimination using Time-over-Threshold (ToT) technique to discriminate between $\gamma$-rays, fast neutrons and the thermal neutrons. The inner wall of the detector is coated with a Lithium-based compound to make the detector sensitive to the thermal neutrons[8]. These neutrons undergo capture by Lithium to produce energetic $\alpha$-particles which result in scintillation inside the detector volume. Corresponding ToT signals are found to be larger than those produced by the fast neutrons. ToT spectra measured with the S670e PHe detector for the neutrons and the $\gamma$-rays emitted from a 252Cf neutron source and a 137Cs $\gamma$-ray source are shown in the Fig. 1. Two different cuts along the ToT axis, shown by the arrows, differentiate between the $\gamma$-rays, fast neutrons and the thermal neutrons as labelled in the figure. One disadvantage of this method of discrimination is considerable overlap between the gamma rays and the fast neutrons or between the fast neutrons and the thermal neutrons. As a result, a clear threshold for discrimination could not be established.
Figure 1: Time over threshold (ToT) plots obtained by exposing the detector separately with a 252Cf gamma ray and neutron source, and a 137Cs gamma ray source. The two arrows mark the low ToT and the high ToT cuts, which separate the regions of response for the gamma rays, the fast neutrons and the thermal neutrons as shown.
Intrinsic detection efficiency of the above-mentioned PHe detector for the neutrons has been extensively studied using neutron Time-of-Flight (ToF) experiment[9] over the energy range of $\sim 0.5-7$ MeV[9]. Maximum efficiency was found to be 6.8% around 1.5 MeV. It varies within the range of $\sim 4-7\%$ over the energy range of 1 - 7 MeV. However, the simulated efficiency values over the same range was found to differ widely, peaking around 11% at $\sim 1$ MeV.
The objectives of the present work are:
  • discrimination of the fast neutrons from the $\gamma$-rays and electrons using fast-slow charge integration method;
  • estimation of threshold for neutron and $\gamma$-ray discrimination;
  • qualitative study of attenuation of the fast neutrons and $\gamma$-rays by high density polyethylene (HDPE or PE);
  • carry out a detailed simulation to a) establish the neutron emission spectrum from a 241Am based radioactive source used in the experiment, b) to understand the interaction of the neutrons and $\gamma$-rays with the PHe media;
  • follow the systematics of energy transfer from the neutrons and the $\gamma$-rays to 4He (nuclear recoil) and electrons (electron recoil) respectively.
@@ 2.1 Instrumentation
A photograph of the experimental set up is shown in the Fig. 2(a). Arktis-made PHe detector and a radioactive source (252Cf source inside a PTFE capsule), placed in front of the detector, are displayed in the photograph. The active medium of the detector is 4He gas at a pressure of 180 bar, enclosed inside a stainless steel tube of 60 mm inner diameter and 600 mm active length. The detector is packaged with SiPM arrays as photon readout located inside the pressurized detector volume. The detector is segmented into three parts along the length of the cylindrical tube. Photons from each segment are read out by an array of 8 SiPMs. Signals from two SiPMs are summed and fed into each output, so that 4 output signal pulses are generated from each segment. Details about the detector, related electronics and their working principles are given in many references available[6, 8].
@@
For our work, the digital electronics readout system was replaced by a multi-channel analog circuit (provided by the manufacturer) for analog signal processing using conventional electronics and a 12-bit current-integrating analog-to-digital converter or QDC (16 channel Phillips Scientific charge ADC, Model 7187). The pulses from the SiPMs are fed to the analog circuit, which consists of a preamplifier and a shaping amplifier with baseline restorer for each SiPM channel. Traces of the output pulses from the amplifier are shown in the Fig. 2(b). The output signals were inverted using high speed pulse transformer (ALT 4532M) to match with the polarity and pulse timing requirements. Out of the four output signals from a segment, one signal was fed to a low threshold discriminator to generate the logic gates for the QDCs required for pulse shape discrimination (PSD) between neutrons and $\gamma$-rays or electrons. The logic gates, viz. the long gate was matched with the timing and the duration of the output signal in order to integrate the whole pulse, while the short gate was chosen by optimization of the PSD signal. The PSD parameter ($P$) is defined as: $P=NQ_{S}/{(Q_{S}+Q_{L})}$ where, $Q_{S}$ and $Q_{L}$ are the charge contents of the pulse within the duration of the short gate and the long gate respectively and $N$ is a scaling factor to convert the ratios to suitable integer values for the plots.
@@ 2.2 Measurements with Radioactive sources
The systematic studies were done using different neutron and $\gamma$-ray sources placed at a certain distance from the detector as shown in the photograph (see Fig. 2(a)). Provision for placement of different absorber materials in between for systematic studies was also made. An unmoderated 252Cf spontaneous fission source (half-life = 2.645 years, spontaneous fission branching ratio = 3.09%[10]) was used for simultaneous detection of neutrons and $\gamma$-rays. The source, with a residual strength of a few kBq, emits fast neutrons with average energy of 2.12 MeV. The spontaneous fission source also emits $\gamma$-rays of various energies ranging within a broad range of $0.1-1$ MeV due to the transitions between the nuclear levels of the residual fission fragments. Since the source had already passed through approximately 8 half-lives since its preparation, the emission rates of $\gamma$-rays compared to that of fast neutrons are higher by at least 2 orders of magnitude. The source was sealed inside a cylindrical PTFE capsule having an opening on one of the flat faces. The aperture was sealed with a 100 $\mu$m thick polyethylene terepthalate (PET) window to prevent the alpha particles, recoiling heavy nuclei including the fission fragments from ejection into the air.
A typical two dimensional (2D) isometric or scatter plot of the PSD parameter $P$ vs. $Q_{L}$ is shown in the Fig. 3 for the 252Cf source placed near the detector. Two distinct bands can be seen in the scatter plot
Figure 3: (a) 2D scatter plot of the PSD parameter $P$ as function of $Q_{L}$ obtained by exposure of 252Cf source. The contours marked around the bands are the banana gates for the neutrons (black) and the $\gamma$-rays (red). (b) A typical signal pulse output of the detector with the short gate (blue) and the long gate (red) is shown.
obtained after 5 hours of exposure to the 252Cf source. The relative configuration of the gates, optimized by adjusting gate widths and delays by looking at $P$, is also shown in Fig. 3b. Optimum delay between the threshold for the up-swing of the pulse and the trigger point of the short gate was found to be around 300 nanosecond for achieving good discrimination.
A 137Cs monoenergetic (662 keV) $\gamma$-ray source, with strength $\sim 30\,{\rm kBq}$, was used for the experiment to identify the $\gamma$-band as distinct from the neutron band. The scatter plot, obtained after exposing the detector with the $\gamma$-ray source, is displayed in Fig. 4, which shows a single band ($\gamma$-band) as expected. Additional confirmation of the neutron band was obtained after systematic attenuation studies made using high density polyethylene (HDPE or PE) absorbers (see Sec. 2.4).
@@
From the 2D scatter plot of Fig. 3a, the neutron and the $\gamma$-bands are found to merge at low $Q_{L}$ values, which qualitatively indicates the low energy threshold for discrimination. Projected spectra of the neutrons and the $\gamma$ bands were taken after selection of the corresponding bands by the contours (popularly termed as banana gates) shown in the 2D scatter plots of Fig. 3a. A comparison of projection of the neutron and the $\gamma$-bands on the $Q_{L}$ axis is shown in Fig. 5. It reveals that the detector response to the $\gamma$-rays, mediated mostly through electron recoil, is quite low compared to that for the neutrons having energies of the same order. It can be seen from the plots that the nuclear recoil spectrum due to the neutrons terminate abruptly at higher channels ($\gtrsim 250$), which is due to the saturation of the pulses of the SiPM signal amplifiers provided by the manufacturer.
@@
2.3 Measurements with 241Am source
A 241Am source (432 years half-life), embedded in a glass matrix and sealed inside a cylindrical stainless steel capsule, was also studied with the detector. The source, having a strength of $\sim 100\,{\rm MBq}$, is originally intended for energy calibration of X-ray detectors. This source was used in the experiment to determine the response of the detector to low energy photons so that a low energy cut-off for electron recoil could be established. The scatter plot of $P$ vs. $Q_{L}$, obtained after 5 hours of exposure, is shown in Fig. 7. A clear neutron band, besides the $\gamma$-band, was observed to appear. This was a bit surprising due to the reason that 1) the source was used so far for energy and relative efficiency calibration of X-ray detectors, and 2) it was never monitored with neutron detectors or dosimeters. The same exposure time (5 hours) for the 241Am and the 252Cf sources to the PHe detector was used.
In order to understand the origin of the neutrons and its associated spectral distribution, a detailed simulation was carried out to understand the production of neutrons from the 241Am radioactive source. This was done using the known geometry and the media surrounding the source material as described in the following paragraph. The simulation study was performed using GEANT4 (G4) simulation toolkit [11], version 10.7.4. A custom physics-list is utilized in our simulation, which is based on the one developed by Mendoza et al.[12]. We will briefly outline the G4 packages used in this physics-list to model various interactions. The G4RadioactiveDecayPhysics package is used for modeling of the radioactive decays, relevant for the 241Am source. It uses the ENSDF database [13] for the various decay parameters including the energy levels of the daughter nuclei. The production of neutrons from ($\alpha,n$) reaction is modeled using the G4ParticleHP package, which is capable of utilizing the ENDF-6 formatted data libraries[14]. The database used for this purpose is the JENDL/AN-2005 dataset [15]. The QGSP_BIC_HP model is used for the hadronic interactions and uses the G4 Neutron Data Library (G4NDL) to implement low energy neutron interactions with high-precision. The EM interactions are modeled with the G4EmStandardPhysics_option4 package.
The geometry to obtain the particle spectra consists of a cylindrical stainless steel source capsule of 6 mm diameter $\times$ 9 mm length, with the 241Am source, embedded in a pyrex glass matrix of 3 mm diameter $\times$ 6 mm thickness, placed inside and sealed with a stainless steel cover. The source capsule is surrounded by a spherical dummy detector, which is made of air. This is placed to track the particles that are being emitted from the source. 241Am is used as the primary $\alpha$-emitter nucleus embedded within the glass matrix. It decays to 237Np by emitting $\alpha$-particles at 5.486 MeV ($\sim 85\%$ decay branch), 5.443 MeV ($\sim 13\%$ decay branch), and the rest at other energies. Np237 daughter nucleus has a half-life of approximately 4 orders of magnitude more than that of 241Am. Therefore, the source is approximated as an alpha emitter whose energy is sampled from the spectrum as obtained from G4. These $\alpha$-particles are allowed to penetrate isotropically through the base material of the source capsule, where they undergo ($\alpha,n$) reaction with the constituent nuclei. Based on the cross-section data available from the database, respective elemental compositions of the materials and the relevant energies of the $\alpha$-particles, we find that the dominant reaction producing neutrons are: 11B($\alpha,n$)14N ($\sim 69\%$) and 23Na($\alpha$, n)26Al($\sim 11\%$), rest are being produced by various other reactions.
Figure 6: Energy spectrum of the neutrons arising from the 241Am source capsule as obtained from G4 simulation.
2.4 Systematic Studies with HDPE Attenuator
In order to confirm that the so-called neutron band is due to the neutrons being detected, attenuation of the band population by the HDPE was studied. For this purpose, 4 layers of HDPE, each being 25 mm thick with a total thickness of 100 mm, were placed between the 241Am source and the PHe detector to attenuate the neutrons. The scatter plot ($P$ vs. $Q_{L}$), obtained over the same exposure time, is shown in Fig. 7.
@@ 2.4 Systematic Studies with HDPE Attenuator
It clearly reveals attenuation of the neutron band. The plots, obtained by gating on the neutron band and projection taken along the $Q_{L}$ axis (see Fig. 8), provide a comparison of the neutron attenuation by the HDPE layers. Low energy neutrons are expected to be attenuated more than those of higher energies, however from the plot, the attenuation factor appears to be fairly independent of energy. This is due to the moderation of the energetic neutrons that enhances counts at low energy and suppresses at higher energy.
Similar projection spectrum for the $\gamma$-band is shown in the Fig. 9, where attenuation of $\gamma$-rays is observed. The $\gamma$-ray spectrum at the lower channels along the $x$-axis get significantly attenuated after passing though the HDPE layers due to the presence of scatterers. However, the higher channels of the spectrum corresponding to larger deposited energy ($E_{dep}$) values get populated as well. For measurements with different number of HDPE layers in between, we observe that after traversing two HDPE layers, the spectrum at the higher channels (i.e region of interest or RoI $>150$) show a rising trend and a bump around 240 channel number till it reaches saturation at the highest channel. There is not much visible relative change for four HDPE layers. However, the spectral population over the RoI diminish after passing through five HDPE layers. This may be interpreted as due to the production of 2.225 MeV $\gamma$-rays from neutron moderation within the HDPE, followed by absorption in HDPE through the $n+p=d+\gamma$ process. A simulation of the $\gamma$-ray interaction with the detector is done to understand the origin, the nature of the $\gamma$-ray spectra and the underlying systematics of the process.
@@
Figure 9: Same as Fig. 8 except that the banana gate is set on the $\gamma$-band. (a) projected spectra with and without the HDPE (PE) layers between the source and the detector. See Sec. 2.4 for the calibration procedure applied in the plot.(b) Comparison of the projected $\gamma$-spectra recorded after the neutrons travel through different numbers of the HDPE layers. The $x$-axis scale on the bottom displays the electron equivalent energy loss $E_{dep}^{ee}$ of neutron. See Sec. 2.4 for its definition.
$\gamma$-rays falling on the PHe detector cause electrons inside the gas to recoil and thus, result in electron recoil signal. This is demonstrated in our experiment (see Fig. 4) with 137Cs standard $\gamma$-ray source as well. The range of $E_{dep}$ values can be seen to merge with the lower range of $E_{dep}$ values for the neutrons as shown in the Fig. 5, where the spectra for the neutrons and the $\gamma$-rays from 252Cf are overlaid. Henceforth, we will designate the energy loss by the electrons as $E_{dep}^{ee}$ to distinguish it from $E_{dep}$ due to the neutrons. The study is more complicated due to the presence of stainless steel housing of the detector which contains high $Z$ elements. Since the $\gamma$-rays are energetic enough to undergo pair production, positrons are also generated, which result in the production of 511 keV $\gamma$-rays. Therefore, the ER events, though largely contain effects due to the electrons, there would be effect due to the positrons as well.
In the simulation procedure, mono-energetic $\gamma$-rays of different energies ($E_{\gamma}$) are allowed to impinge upon the detector. The $E_{dep}^{ee}$ spectrum of the recoiling electrons is shown in Fig. 10. A peak begins to appear at $E_{dep}^{ee}\sim 280$ keV as the energy of the incident $\gamma$-ray is increased. The $\gamma$-ray energy threshold for appearance of the peak is found to be around 650 keV. Therefore, the 847 keV $\gamma$-ray, which arises from the ${}^{56}{\rm Fe}(n,n^{\prime}\gamma)$ reaction due to the presence of SS housing, would also give rise to the $\sim 280$ keV peak, when no HDPE slab is placed. This explains the very few counts observed in the higher channel numbers in the absence of HDPE attenuator in Fig. 9 (black line). A detailed event-by-event tracking reveals that this number corresponds to the most probable $E_{dep}^{ee}$ of the recoiling electrons given the geometry (diameter) of the detector.
Figure 10: $E_{dep}^{ee}$ spectrum of mono-energetic $\gamma$-rays
The peak-like feature in the Fig. 10 arises due to the fact that the recoiling electrons do not lose their entire energy within the detector. This can be seen more explicitly in Fig. 11, where the $E_{dep}^{ee}$ is plotted against their recoil kinetic energy (KE) for mono-energetic incident $\gamma$-rays of 1 MeV and 2.225 MeV. The kinetic energy of the recoiling electrons increases with incident energy of the $\gamma$-rays as expected. The sharp right edge in the plots is the Compton edge. It is evident from Fig. 11 that a significant fraction of the electrons with higher recoil energies lose only $\sim 280$ keV of their energy within the detector. Therefore, we can conclude that the peak observed at higher channel numbers in the Fig. 8 corresponds to $\sim 280$ keV of $E_{dep}^{ee}$. To further confirm this geometric effect, the simulation was repeated by changing the diameter of the fiducial volume of the detector to 30 mm and 90 mm. Corresponding $E_{dep}^{ee}$ spectra are plotted in the Fig. 12. A clear shift in the peak can be seen, which is at $\sim 130$ keV for 30 mm diameter and $\sim 450$ keV for 90 mm.
[…]
Figure 11: $E_{dep}^{ee}$ vs. recoil kinetic energy of ER events from mono-energetic $\gamma$ sources for (a) 1.0 MeV and (b) 2.225 MeV
Based on the experimental and the simulation results on the interaction of the neutrons with HDPE and $\gamma$-ray response of the PHe detector, we can conclude that: 1) the build-up observed in the higher channel number is due to the production of 2.225 MeV $\gamma$-rays from $n+p=d+\gamma$ process 2) the $E_{dep}^{ee}\sim 280$ keV peak arises due to the most probable value of energy loss by the recoiling electrons within the confined geometry of the PHe detector. This is also seen as a rising bump in the RoI at the higher channel numbers ($\sim 240$) in the experiment (see Fig. 9), apart from the spectral feature at the lower channel numbers ( $\leq\,150$) which has been demonstrated as arising from accompanying low energy $\gamma$-rays. Based on the spectral profile of the $E_{dep}^{ee}$ spectra, we fitted the peak with logistic function after subtracting an exponential background for the $E_{dep}^{ee}$ spectra shown in the Fig. 10, covering 662 keV to 2.225 MeV $\gamma$-rays. The average $E_{dep}^{ee}$ value for the peak was found to be: $282\pm 10$ keV. Considering this and the fact that the $x$-axis of the plots of the Fig. 9 scale as $E_{dep}^{ee}$, a first order calibration with calibration constant of 1.175 keV/channel on the $x$-axis is applied.
Figure 12: $E_{dep}^{ee}$ spectra with 30 mm (solid lines), 60 mm (dot-dashed lines) and 90 mm (dotted lines) diameter of the active volume of the detector.
2.5 Neutron Energy Calibration based on 252Cf Spectral Data